Paolo and JOse drove in opposite directions. After three hours they were 63 km apart. Paola drove 3 km faster tha Jose. How fast did Paola drive?
Answer by shree840(260)

let joso drive at the rate of x miles per hour.
paola drove at x+3 rate
in three hours joso drove 3x miles
paola drove 3(x+3) miles=3x+9
total miles =3x+3x+9=63
6x+9=63
6x=54
x=9
joso 9 miles per hpur
Paola 12 miles per hour